Light generator crate guidance
This webpage does not aim at providing a complete technical documentation of the DIRC light generator crate: such information can be found in the DIRC internal note 118. It rather presents several tests which can be tried to understand the origin of a calibration problem involving the light generator crate. They are based on the investigations done at the end of the October 2005 shutdown to understand a calibration problem located in Sector #10 -- see this thread for more details on that particular problem.
Before investigated any problem, having a printed copy of note 118 is certainly a good idea. In particular, the DIRC calibration schematic drawn in page 2 can be very useful to find what a particular test proves and what it doesn't.
When a DIRC calibration is taken, the front-end electronics send triggers to the LEDs which are in the light generator crate and which in turn send light pulses inside the SOB. Then, for each PMT, the distribution of the recorded hit arrival times is fitted by a complex function to extract its timing correction. Other sections of this manual show the way to take a standalone calibration and to check the results of a DIRC calibration (either standalone or global, i.e. run from the ORC interface).
Short description of the light generator crate components
The two pictures below show the front and rear sides of the DIRC light generator crate which is located in the E.H., above the HV crates powering sectors #0 to #5.
Front side of the DIRC light generator crate.
On the front side, 6 different components are visible.
- 4 'Boosters', labelled 1 to 4.
A booster contains 3 blue diodes which send light in 3 different sectors through the optical fibers which are plugged on the board front side. In the back side, there are:
We also have one spare booster (labelled '5') which was tested OK in Saclay -- actually this simply means that the diodes were working fine as there is no light generator crate in Saclay: the spare is here at SLAC!
- 3 fibers which send light from the diodes to the PMT located inside the crate (and used to test the LEDs one by one, see note 118 for details);
- 1 big Cannon outlet which brings power and triggers to the boosters.
- 1 'Internal Trigger Module for the boosters'.
This is the largest component in the center of the crate; On the picture, its digital screen shows the voltage applied to the LEDs and its main mode switch is on 'EXTERNAL' (top position): this means that the LED pulses are triggered by the DIRC front-end electronics -- which is the normal configuration when BaBar is running. When the switch is on the other position ('INTERNAL'), the diodes can be internally triggered. This configuration is interesting for tests as it allows one to take data in standalone mode while the LEDs are flashing.
The 12 smaller switches allow one to turn ON (switch to the left) or OFF (switch to the right) any of the 12 diodes. When a diode is ON, a green light should be visible on the left of the corresponding switch -- in October 2005, green light #7 was found dead.
- 1 'Converter Control Module'.
This is the leftmost blue module on the picture above. Its main interests are twofold:
- to run ON or OFF the HV of the PMT located inside the crate;
- to select the frequency of the internal pulse generator -- 10 kHz (the max. frequency) should be the best choice for tests on the DIRC .
Rear side of the DIRC light generator crate.
The above picture shows the rear side of the light generator crate. On the left side, one can see the black main power switch which is used to turn ON/OFF the whole crate. The 12 white cables bring information from the DIRC front-end crates -- there are also 2 spare cables not connected which could be used in case of a major problem with one of the wires currently used. Obviously, they would also need to be connected to one DIRC sector on the other side!
Testing the calibration procedure
This part presents some tests which could be useful to understand a problem of unknown origin. Feel free to add more tests of our own if you find some and to change the conclusions we draw from them if you find them inaccurate.
- Swapping 2 optical fibers on the booster front panels before taking a standalone calibration. In that case, the LED target sectors are exchanged accordingly but each diode is still triggered by the same front-end crate as in the original configuration.
- Swapping two CANNON outlets at the back of two boosters: the diodes get steered by different crates but they still emit in the same sectors.
- The same effect can be achieved by swapping 2 cables at the rear side of the light generator crate.
- Using the crate internal trigger to take standalone data with the suitable set of diodes sending pulses in the SOB at high rate is a good test. One could for instance see how the DIRC is illuminated when the 12 diodes are working together or compare the lighting of two diodes, turned ON one after the other in two different runs.
- If the problem is traced back to the connection between one front-end sector and the light generator, checking the cable end connector (by plugging in and out etc.) can be a good idea as well. Indeed, that's the way the sector 10 calibration problem was fixed in October 2005!
Pending and done tests
During spring 2005 Jerry and Nicolas tested the PMT in the spare light generator crate: after a few discussion with Saclay people (mainly Philippe Bourgeois), it was found to be working.
On the other hand, we didn't succeed in taking a standalone calibration when we used the spare booster during the October'05 shutdown: the calibration stopped almost immediately. Due to lack of time (and to the fact that we fixed the problem without the need of the spare booster!), the reason for the failure was not understood. Yet, the problem is likely to come from the current settings of the LEDs. For tests in Saclay they may have been tuned to maximum power whereas an emission lower by an order of magnitude is needed for the DIRC. Confirming or not this hypothesis would require to look at the internal PMT response when it is lighted by these diodes.
A few plots produced during the October 2005 shutdown
Web Page by Nicolas Arnaud
Last updated on November 7th, 2005.